one number is three times another number and their sum is 164.find the number?
: one number is three times another number and their sum is 164.find the number?

Answer by ptaylor(1170)

Let x=the other other number
Then 3x=the first number
Now we are told that:
x+3x=164 or
4x=164 divide both sides by 4
x=41----------------------------the other number
3x=3*41=123------------------------the first number
CK
41+123=164
164=164
Another way:
Let x=the first number
And let y=the second number
x+y=164-----------------------eq1
x=3y-----------------------------eq2
Substitute x=3y from eq2 into eq1 and we get:
3y+y=164
4y=164
y=41-------------the second number
substitute y=41 into eq2
x=3y=3*41=123-------the first number

Hope this helps---ptaylor